OpenStudy (anonymous):

You are making poster board boundary markers in the shape of square pyramids for a relay race. If you are including a ten -inch square as the bottom and you want the slant height to be twelve inches, how much poster board do you need?

OpenStudy (anonymous):

three hundred forty square inches three hundred sixty square inches three hundred eighty-four square inches four hundred square inches

OpenStudy (anonymous):

@LegendarySadist help

OpenStudy (anonymous):

In other words, we're finding the surface area of a square pyramid, right?

OpenStudy (anonymous):

here i think not sure but here is an image that may be that or not if it is not then yea your right sorry get confuse with shapeshttp ://learn.flvs.net/webdav/assessment_images/educator_math2/v9/07_TB_39_240.jpg

OpenStudy (anonymous):

I got a cookie error for the image /:

OpenStudy (anonymous):

really okay give me a second

OpenStudy (anonymous):

this is from a different test question but just to give an idea if that the shape we think it is

OpenStudy (anonymous):

Well anyways. With the info we have, we can put it into the formula \[\huge (a^{2})+(4*\frac{1}{2}*sh*a)\] with a being the side of the square and sh being the slant height

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

yea just calculating and i also went to make coffee for my mom :D

OpenStudy (anonymous):

250 is that it

OpenStudy (anonymous):

What did you use for a and what did you use for sh?

OpenStudy (anonymous):

10 for the area and 12 for the slant height

OpenStudy (anonymous):

So \[\large 10^{2}=~?\]

OpenStudy (anonymous):

20

OpenStudy (anonymous):

100

OpenStudy (anonymous):

i am dummy its 100

OpenStudy (anonymous):

Ok yeah, 100. So \[\large 4*\frac{1}{2}*12*10=~?\]

OpenStudy (anonymous):

240 + 100 = 340

OpenStudy (anonymous):

Right! So 340 is gonna be our answer :)

OpenStudy (anonymous):

yay thank you so much i keep that formula in mind

OpenStudy (anonymous):

\[\huge \color{aqua}N\color{fuchsia}o \space \color{lime}P \color{orange}r \color{blue}o \color{maroon}b \color{red}l \color{olive}e \color{purple}m \ddot\smile \]

OpenStudy (anonymous):

cool how did you do that

OpenStudy (anonymous):

Latex writing, see the code-like writing in my bio?

OpenStudy (anonymous):

cool okay thanks for the help